OpenStudy (anonymous):

Graph the equation using the slope and the y-intercept y=4/3x + 7

OpenStudy (saifoo.khan):

i will give the same answer!

OpenStudy (saifoo.khan):

when x intercrpts y = 0 whn y intercpts x = 0

OpenStudy (anonymous):

I think you asked this already.. Start by finding the y-intercept (plug in 0 for x). Then go up 4 and right 3 (according to the slope) Then draw the line between those points.

OpenStudy (saifoo.khan):

by loooking @ the equation, u will C 7. this is the y intercpt.

OpenStudy (saifoo.khan):

now find the x intercpt.. y=4/3x + 7 0 =4/3x + 7 -7*3 ---- = x 4 x = -21/4

OpenStudy (anonymous):

that's a yucky x-intercept. Why not just move up 4 and right 3 for a second point.

OpenStudy (anonymous):

very confused

OpenStudy (anonymous):

1) Start by finding the y-intercept (plug in 0 for x). 2) Then go up 4 and right 3 (according to the slope) to find a second point 3) Then draw the line between those points. Which part confuses you, and why?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

You solved it?

OpenStudy (anonymous):

no

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

We are not mind readers. If you are confused, please explain what confuses you. Otherwise we just have to guess and will probably be wrong. We cannot help you if you cannot explain what is the source of your confusion.

OpenStudy (anonymous):

I drew you a picture on your other question, but I'll attach it here to. Might help.